The time for a major exam to be completed is normally distributed with an average of 55 minutes and a standard deviation of 9 minutes. If 92% of the students completed the exam, when should the test be terminated?


1
Expert's answer
2021-11-16T13:59:59-0500

"P(Z<z)=0.92 \to z=1.405."

"\frac{x-55}{9}=1.405."

"x=55+9*1.405\approx 68 min."